Please, I need help with the next step, something is not right.
.05(113 – 2y) + 8 + .10(2y) = $10.45

Found 2 solutions by Alan3354, checkley77:
Answer by Alan3354(69443)

.05(113 – 2y) + 8 + .10(2y) = $10.45
-----------
Multiply by 100
5*(113 - 2y) + 800 + 10*2y = 1045
565 - 10y + 800 +20y = 1045
1365 + 10y = 1045
10y = - 325
y = -32.5 cents

Answer by checkley77(12844)

.05(113 – 2y) + 8 + .10(2y) = $10.45
5.65-.1Y+8+.2Y=$10.45
.1Y=10.45-5.65-8
.1Y=-3.2
Y=-3.2/.1
Y=-32 ANSWER.
PROOF:
.05(113-2*-32)+8+.1(2*-32)=10.45
.05(113+64)+8+.1*-64=10.65
.05*177+8-6.4=10.65
8.85+8-6.4=10.65
10.45=10.45